# Quaternion-Domain Super MDS for Robust 3D Localization

### v1

@@ V-C1 Iterative Convergence Analysis
Before comparing the performance of the algorithms, we first examine the convergence behavior of the iterative QD-MRC-SMDS. The matrix $\left[\bm{K}_{2}^{\mathsf{H}}\ \bm{K}_{3}^{\mathsf{H}}\right]$ exhibits a rank-$1$ structure, and (49) is structurally equivalent to the power iteration method. As a result, the direction of $\bm{\nu}_{\mathrm{AT}}$ is expected to converge within a single iteration. Furthermore, since the known vector $\bm{\nu}_{\mathrm{AA}}$ appears explicitly in (49), the scaling of $\bm{\nu}_{\mathrm{AT}}$ is also preserved. Thus, both the direction and scale of $\bm{\nu}_{\mathrm{AT}}$ are anticipated to converge after just one iteration.
Fig. 6 shows the average estimation error of the iterative QD-MRC-SMDS algorithm as a function of the number of iterations $\tau$, under distance error conditions of $\sigma_{\mathrm{d}}=2$ [m] and $4$ [m]. The simulation setup follows the conditions described Section IV-C. As theoretically anticipated, the algorithm converges within a single iteration. This confirms that setting $\tau_{\mathrm{max}}=1$ is sufficient, and the additional computational cost introduced by the iterative procedure corresponds to computing (49) only once.
@@ V-C2 Simulation Results
Under the same simulation conditions as in Figs. 3 and 4, the localization accuracy of the QD-MRC-SMDS and iterative QD-MRC-SMDS algorithms is compared.
Fig. 7 shows the localization performance of each method in Scenario I, with the QD-SMDS algorithm included as reference (black curve). Notably, both QD-MRC-SMDS and its iterative variant achieve high localization accuracy without relying on SVD-based noise suppression. Even the lowest-complexity QD-MRC-SMDS incurs only a modest performance degradation compared to QD-SMDS. Moreover, its ability to maintain high accuracy under large angle errors indicates that QD-MRC-SMDS inherits the robustness to angular uncertainties observed in QD-SMDS. When compared to the SMDS performance shown in Fig. 3, QD-MRC-SMDS demonstrates superior localization accuracy, especially severe angular error conditions (i.e., $\epsilon\geq 30^{\circ}$), underscoring its robustness and computational efficiency. Furthermore, the iterative QD-MRC-SMDS algorithm achieves performance nearly equivalent to that of QD-SMDS, suggesting that the full potential of the GEK matrix can be exploited without requiring low-rank truncation via SVD.
Fig. 8 shows the localization performance of each method in Scenario II. While the overall trend is similar to that observed in Scenario I, the performance gap between the iterative QD-MRC-SMDS and QD-SMDS increases with larger angular errors. This can be attributed primarily to the greater degradation in the accuracy of the GEK matrix caused by increased measurement errors in the additional angular parameters of the azimuth and elevation angles. In this case, it appears that noise suppression in the SVD-based approach works slightly better than in the iterative one. Nevertheless, the performance gap remains minimal—even at the largest angular error ($\epsilon=50^{\circ}$), the estimation error difference is within $0.05$ [m].
@@
Based on these results, system designers can select the most appropriate algorithm by balancing the computational complexity against localization accuracy. For latency-critical applications, the low-complexity QD-MRC-SMDS offers the most practical choice. Conversely, when computational resources are sufficient and maximum accuracy is required—even under large angular errors—QD-SMDS becomes the preferred option. Lastly, the iterative QD-MRC-SMDS achieves an excellent trade-off between accuracy and complexity, making it a strong candidate for a wide range of practical deployment scenarios.
VI Conclusion

### v2

@@ V-C1 Iterative Convergence Analysis
Figure 12: Empirical CDFs of QD-SMDS, QD-MRC-SMDS, and iterative QD-MRC-SMDS algorithms in Scenario I.
Before comparing the performance of the proposed algorithms, this section investigates the convergence behavior of iterative QD-MRC-SMDS. By focusing on the matrix-vector multiplication structure in the numerator of (50), the update rule can be interpreted as a process equivalent to the power iteration method [40]. Whereas the conventional power iteration updates all vector components simultaneously, the proposed method keeps the known component $\bm{\nu}_{\mathrm{AA}}$ fixed and updates only the unknown component $\bm{\nu}_{\mathrm{AT}}$ through multiplication with the matrix $\bm{K}$. Accordingly, the proposed update rule can be viewed as a power iteration formulated within a block coordinate update framework. As a result, $\bm{\nu}_{\mathrm{AT}}$ is expected to converge through the iterations toward the direction aligned with the singular vector corresponding to the largest singular value of the kernel matrix $\begin{bmatrix}\bm{K}_{2}^{\mathsf{H}}\ \bm{K}_{3}^{\mathsf{H}}\end{bmatrix}$.
On the other hand, since the convergence rate depends on the choice of the initial vector, it is generally difficult to derive a theoretically exact number of iterations required for convergence. Hence, the performance variation with respect to the number of iterations is evaluated through numerical simulations in order to characterize the practical convergence behavior of the proposed algorithm.
Fig. 10 shows the average estimation error of the iterative QD-MRC-SMDS algorithm as a function of the number of iterations under distance error conditions of $\sigma_{\mathrm{d}}=2$ m and $4$ m. All other simulation parameters are identical to those used in Section IV-C. As observed from the figure, the performance converges within a single iteration under both conditions. It is well known that the power iteration method converges more rapidly toward the dominant direction when the rank-one component is dominant. Since the block-structured kernel matrix $\begin{bmatrix}\bm{K}_{2}^{\mathsf{H}}\ \bm{K}_{3}^{\mathsf{H}}\end{bmatrix}$ exhibits a structure close to that of a rank-one matrix, this behavior is in good agreement with that intuition. This confirms that setting $\tau_{\mathrm{max}}=1$ is sufficient, and the additional computational cost introduced by the iterative procedure corresponds to evaluating (50) only once.
@@ V-C2 Simulation Results
Figure 14: Comparison of empirical CDF between the QD-SMDS, QD-MRC-SMDS, and iterative QD-MRC-SMDS algorithms in Scenario II.
In this subsection, we compare the localization accuracies of the QD-MRC-SMDS and iterative QD-MRC-SMDS algorithms under the same simulation conditions as in Figs. 4 and 6 and under the two distinct scenarios.
Comparisons under Scenario 1: Figs. 11 and 12 show the average estimation error and the empirical CDF of both methods in Scenario I, with the SMDS algorithm included as reference (black curve). Notably, both QD-MRC-SMDS and its iterative variant achieve high localization accuracy without relying on SVD-based noise suppression. Even the lowest-complexity QD-MRC-SMDS incurs only a modest performance degradation compared to QD-SMDS. Moreover, its ability to maintain high accuracy under large angle errors indicates that QD-MRC-SMDS inherits the robustness to angular uncertainties observed in QD-SMDS. When compared to the SMDS performance shown in Fig. 4, QD-MRC-SMDS demonstrates superior localization accuracy, especially severe angular error conditions (i.e., $\epsilon\geq 30^{\circ}$), underscoring its robustness and computational efficiency. Furthermore, the iterative QD-MRC-SMDS algorithm achieves performance nearly equivalent to that of QD-SMDS, suggesting that the full potential of the GEK matrix can be exploited without requiring low-rank truncation via SVD.
Comparisons under Scenario 2: Next, Figs. 13 and 14 show the average estimation error and the empirical CDF of each method in Scenario II, respectively. While the overall trend is similar to that observed in Scenario I, the performance gap between the iterative QD-MRC-SMDS and QD-SMDS increases with larger angular errors.
This can be attributed primarily to the greater degradation in the accuracy of the GEK matrix caused by increased measurement errors in the additional angular parameters of the azimuth and elevation angles. In this case, it appears that noise suppression in the SVD-based approach works slightly better than in the iterative one. Nevertheless, the performance gap remains minimal—even at the largest angular error ($\epsilon=50^{\circ}$), the average estimation error difference is within $0.05$ m.
VI Computation Effort Comparisons and Discussions
VI-A FLOP Count of SMDS-based Approaches
To enable a fair comparison of the computational effort among the SMDS variants presented in this paper, the total number of FLOPs required by each algorithm is reported.
A detailed derivation of the FLOP count for each algorithm is provided in the Appendix and summarized in Table II, which also includes the FLOP count for the configuration used in the previous simulations.
For the given $N_{A}$ and $N_{\mathrm{T}}$, QD-MRC-SMDS is observed to reduce the computational effort by three orders of magnitude compared with QD-SMDS, because it only requires the computation in (48). Iterative QD-MRC-SMDS slightly more expensive than QD-MRC-SMDS, followed by SMDS and, finally, QD-SMDS. Notably, both QD-MRC-SMDS and its iterative variant exhibit lower computational cost than SMDS, even though the computations are performed in the quaternion domain.
The advantages of switching from SMDS to its computationally efficient quaternion-domain variants is evident when combining the analysis of the overall localization performance with the computational effort in Table II. These variants provide improved localization performance for larger angular errors while requiring lower computational effort.
@@ VI-A FLOP Count of SMDS-based Approaches
Based on these results, system designers can select the most appropriate algorithm by balancing the computational effort against localization accuracy. For latency-critical applications, the low-complexity QD-MRC-SMDS represents the most practical choice.
Conversely, when computational resources are sufficient and maximum accuracy is required—even under large angular errors—QD-SMDS becomes the preferred option. Finally, iterative QD-MRC-SMDS achieves an excellent trade-off between accuracy and complexity, making it a strong candidate for a wide range of practical deployment scenarios.
TABLE II: FLOP count formulas for all SMDS variations. The total FLOP count for the configuration used in the reported simulations is also provided.
Algorithm FLOP count formula FLOP count for $N_{A}=5,N_{\mathrm{T}}=15$
SMDS $3M^{2}+\frac{20}{3}M^{3}+3(M+1)$ $3(N_{\mathrm{A}}+N_{\mathrm{T}})(2(N_{\mathrm{A}}+M)-1) +$$54N_{\mathrm{A}}+21N_{\mathrm{T}}+397.$ $\approx 4.13\times 10^{6}$
QD-SMDS $12M^{2}+\frac{640}{3}M^{3}+4M+1$ $3(N_{A}+N_{\mathrm{T}})(2(M+N_{A})-1) +$$54N_{\mathrm{A}}+21N_{\mathrm{T}}+397.$ $\approx 1.31\times 10^{8}$
QD-MRC-SMDS $12M^{2}+16N_{A}N_{\mathrm{T}}\left(N_{A}(N_{A}-1)-1\right)$ $8(N_{A}(N_{A}-1)/2)-1+8N_{A}N_{\mathrm{T}} +$$N_{\mathrm{T}}(8N_{A}-4)$ $\approx 1.11\times 10^{5}$
Iterative QD-MRC-SMDS $12M^{2}+16N_{A}N_{\mathrm{T}}\left(N_{A}(N_{A}-1)-1\right)$ $8(N_{A}(N_{A}-1)/2)-1+8N_{A}N_{\mathrm{T}} +$$N_{\mathrm{T}}(8N_{A}-4)+16N_{A}N_{\mathrm{T}}(2M-1) +$$8M-1+4N_{A}N_{\mathrm{T}}$ $\approx 3.14\times 10^{5}$
VI-B Complexity Comparisons with Other Localization Methods
To place our QD-SMDS approaches in context, we compare the computational complexity of our methods with other localization algorithms. Among convex optimization approaches, most semidefinite programming (SDP)-based methods tie the complexity primarily to the number of targets $N_{\mathrm{T}}$. For fixed dimension and typical SDP solvers, the complexity grows polynomially but very steeply with $N_{\mathrm{T}}$, with a worst case bound of $O(N_{\mathrm{T}}^{6})$ [46]. In contrast, the proposed QD-SMDS approaches scale quadratically in $N_{\mathrm{T}}$, which makes them more suitable than SDP-based solvers for large-scale networks.
A comparison with Bayesian approaches is more nuanced, since their complexity typically depends on additional framework-specific parameters. In the method proposed in [13], the hybrid Bayesian message-passing algorithm scales linearly with the number of targets, with a per-iteration cost on the order of $O(N_{A}N_{\mathrm{T}}MN)$, where $M$ and $N$ denote the transmitted and drawn sample sizes, respectively. Similarly, for other Bayesian methods [14, 15], the per-iteration cost is linear in $N_{\mathrm{T}}$. However, unlike Bayesian or message-passing localization algorithms–which require multiple iterations and repeated processing of all anchor–target measurements–QD-SMDS provides a single-pass, closed-form computation. In particular, the QD-MRC-SMDS variant scales linearly in $N_{\mathrm{T}}$ without requiring iterative refinement.
A third class of competing techniques consists of machine-learning-based localization methods [47, 48, 49, 50, 51], which learn a mapping from measured features to positions. This line of indoor positioning research employs deep neural networks using channel state information (CSI) fingerprints and explicitly separates a computationally intensive offline training phase from a lightweight online localization phase consisting of a single forward pass through the network. Thus, while learning-based methods trade substantial iterative offline complexity and data-collection effort for very low online cost, QD-SMDS offers a non-iterative, training-free alternative with explicit and predictable complexity for any given $N_{A}$ and $N_{\mathrm{T}}$.
VI-C Implementation Feasibility and Potential Deployment Challenges
Although the evaluation of QD-SMDS in this paper is conducted through controlled simulations, we emphasize that real-life implementation of the method can be easily carried out since the technique is an MDS-based approach [52], for which explicitly experimental results can be found in related literature [53, 54, 55, 56, 57].
To elaborate further, the method is compatible with practical range- and angle-based localization systems. The distance measurements required by QD-SMDS can be obtained from widely used technologies such as ultra-wideband (UWB) time of arrival (ToA)/time difference of arrival (TDoA) devices or received signal strength indicator (RSSI)-based ranging modules. Prior work using these technologies has demonstrated that such systems can provide sufficiently accurate range estimates under realistic conditions [41, 42].
In addition to range measurements, QD-SMDS utilizes azimuth and elevation angle estimates to form the GEK matrix. Such angle information can be obtained using planar antenna arrays or coprime planar arrays capable of two-dimensional DoA estimation. Practical array-processing techniques—including polynomial-rooting approaches and subspace-based methods such as MUSIC and ESPRIT—enable the extraction of both azimuth and elevation angles from compact antenna configurations [34, 43, 44].
However, the deployment of distance- and angle-capable antenna arrays introduces additional considerations. Accurate DoA estimation requires careful antenna calibration, adequate signal-to-noise ratio, and mitigation of multipath propagation, all of which may otherwise bias angle measurements. In addition, clock synchronization for ToA/TDoA systems, antenna calibration for UWB and array-based platforms, and anchor placement geometry are known to influence the quality of the collected range measurements [45].
Consistent with most of the localization literature, the present work focuses on the algorithmic contribution and its performance under simulated situations. A dedicated real-world deployment and measurement campaign therefore represents a promising direction for future work.
VII Conclusion
@@
-A FLOP Count Calculations
-A1 SMDS
The SMDS algorithm starts by building the GEK matrix in (22). This computation yields a matrix with entries $\cos(\alpha_{mn})d_{m}d_{n}$ at row $m$ and column $n$, requiring two multiplications per matrix entry. Additionally, we also need to evaluate the cosines of the measured angles, and the number of FLOPs for such an operation is hardware-dependent. For simplicity, we will assume that one cosine evaluation is equivalent to 1 FLOP, yielding
[…]
In principle, the SVD can be computed using a truncated method that returns only the three dominant eigenpairs [32]. However, the corresponding FLOP count depends on the number of iterations required for convergence. Therefore, to avoid iteration-dependent FLOP counts, we assume that a full SVD is used, for which the number of FLOP s is approximately [40]
[…]
The next operation in (23) involves three square-root operations and multiplying each column by the corresponding singular value. For simplicity, we assume that a square-root operation is equivalent to one FLOP, yielding
[…]
The next operation is (24), where the pseudoinverse
[…]
is computed. Since this pseudoinverse needs to be computed only once and reused across simulations, we assume it is already available and do not include its computation in the total FLOP count. Thus, computing (24) amounts to
[…]
In the final step, the estimated coordinate system is aligned with the true coordinate system using the Procrustes transformation [31]. The alignment consists of three operations: translation, scaling, and rotation.
First, both coordinate systems are translated such that their centroids coincide with the origin. For the true anchor-node coordinate matrix $\bm{X}_{\mathrm{A}}\in\mathbb{R}^{N_{\mathrm{A}}\times 3}$, the centroid is given by
$\bar{\bm{X}}_{\mathrm{A}}=\frac{1}{N_{\mathrm{A}}}\sum_{k=1}^{N_{\mathrm{A}}}\bm{X}_{\mathrm{A},k},$ (60)
where $\bm{X}_{\mathrm{A},k}$ denotes the coordinates of the $k$-th anchor node. The centered coordinates are then obtained as
$\bm{X}_{\mathrm{A},k}^{(\mathrm{O})}=\bm{X}_{\mathrm{A},k}-\bar{\bm{X}}_{\mathrm{A}}.$ (61)
The same centering procedure is applied to the estimated anchor coordinates.
Next, the centered matrix $\bm{X}_{\mathrm{A}}^{(\mathrm{O})}$ is normalized using its Frobenius norm,
$\bm{X}_{\mathrm{A}}^{(\mathrm{norm})}=\frac{\bm{X}_{\mathrm{A}}^{(\mathrm{O})}}{\|\bm{X}_{\mathrm{A}}^{(\mathrm{O})}\|_{\mathrm{F}}},$ (62)
and analogously for the estimated coordinate matrix.
The optimal rotation is obtained by first computing the cross-covariance matrix
$\bm{S}=\left(\bm{X}_{\mathrm{A}}^{(\mathrm{norm})}\right)^{\mathsf{T}}\hat{\bm{X}}_{\mathrm{A}}^{(\mathrm{norm})},$ (63)
followed by its singular value decomposition
$\bm{S}=\bm{U}\bm{\Sigma}\bm{V}^{\mathsf{T}}.$ (64)
The rotation matrix is then given by
$\bm{R}=\bm{V}\bm{U}^{\mathsf{T}}.$ (65)
Finally, the aligned estimate is obtained as
$\hat{\bm{X}}_{\mathrm{aligned}}=s\,\hat{\bm{X}}\bm{R}+\bar{\bm{X}}_{\mathrm{A}},$ (66)
where the scaling factor is
$s=\frac{\|\bm{X}_{\mathrm{A}}^{(\mathrm{O})}\|_{\mathrm{F}}}{\|\hat{\bm{X}}_{\mathrm{A}}^{(\mathrm{O})}\|_{\mathrm{F}}}.$ (67)
Since the true anchor-node coordinates are static throughout the simulations, quantities depending solely on $\bm{X}_{\mathrm{A}}$ are computed once and are therefore excluded from the reported floating-point operation (FLOP) counts.
The computational complexity of the Procrustes alignment is summarized as follows:
  • Normalization of $\hat{\bm{X}}_{\mathrm{A}}$:
    $15N_{\mathrm{A}}.$ (68)
  • Computation of the rotation matrix:
    $18N_{\mathrm{A}}+396.$ (69)
  • Final alignment transformation:
    $21N_{\mathrm{A}}+21N_{\mathrm{T}}+1.$ (70)
Combining all contributions yields
[…]
Adding everything together yields
FLOPs SMDS $\approx 3M^{2}+\frac{20}{3}M^{3}+3(M+1)$
$+3(N_{\mathrm{A}}+N_{\mathrm{T}})(2(N_{\mathrm{A}}+M)-1)$
$+54N_{\mathrm{A}}+21N_{\mathrm{T}}+397.$ (72)
where $M=N_{A}(N_{A}-1)/2+N_{A}N_{T}$.
-A2 QD-SMDS
For QD-SMDS (and the other two quaternion-domain variants), the quaternion GEK matrix in (32) must first be constructed. By inspection, this costs four times as much as constructing the real GEK matrix in (22), yielding
[…]
The next operation is QSVD, which amounts to performing an SVD on a $2M\times 2M$ complex matrix. On average, a complex arithmetic operation is four times more expensive than a real one. Thus, we approximate the FLOP count for QSVD as
[…]
Next, the quaternion edge vector is estimated via (33), which involves one square-root operation and $M$ quaternion-scalar multiplications, i.e.,
[…]
Finally, the quaternion edge vector is converted to the real domain, after which the Procrustes transformation and (24) are applied as in the SMDS case. This yields a total FLOP count of
FLOP s for QD-SMDS $=12M^{2}+\frac{640}{3}M^{3}+4M+1$
$+3(N_{A}+N_{\mathrm{T}})(2(M+N_{A})-1)$
$+54N_{\mathrm{A}}+21N_{\mathrm{T}}+397.$ (76)
-A3 QD-MRC-SMDS
After constructing the quaternion GEK matrix, QD-MRC-SMDS reduces to computing (47). First, note that, due to the structure of
$\bm{B}_{\mathrm{AA}}$ $\triangleq\bm{I}_{N_{\mathrm{A}}}\otimes\bm{1}_{N_{\mathrm{T}}\times 1}$
$=\begin{bmatrix}\bm{1}_{N_{\mathrm{T}}\times 1}&\bm{0}_{N_{\mathrm{T}}\times 1}&\cdots&\bm{0}_{N_{\mathrm{T}}\times 1}\\ \bm{0}_{N_{\mathrm{T}}\times 1}&\bm{1}_{N_{\mathrm{T}}\times 1}&\cdots&\bm{0}_{N_{\mathrm{T}}\times 1}\\ \vdots&\vdots&\ddots&\vdots\\ \bm{0}_{N_{\mathrm{T}}\times 1}&\bm{0}_{N_{\mathrm{T}}\times 1}&\cdots&\bm{1}_{N_{\mathrm{T}}\times 1}\end{bmatrix}\in\mathbb{R}^{N_{\mathrm{A}}N_{\mathrm{T}}\times N_{A}},$ (77)
$\bm{B}_{\mathrm{AT}}\triangleq\bm{1}_{N_{\mathrm{A}}\times 1}\otimes\bm{I}_{N_{\mathrm{T}}}=\begin{bmatrix}\bm{I}_{N_{\mathrm{T}}}\\ \bm{I}_{N_{\mathrm{T}}}\\ \vdots\\ \bm{I}_{N_{\mathrm{T}}}\end{bmatrix}\in\mathbb{R}^{N_{\mathrm{A}}N_{\mathrm{T}}\times N_{\mathrm{T}}},$ (78)
the product $\bm{B}_{\mathrm{AA}}\bm{\chi}_{\mathrm{A}}$ simply repeats each entry of the quaternion vector $\bm{\chi}_{\mathrm{A}}\in\mathbb{H}^{N_{A}\times 1}$, $N_{\mathrm{T}}$ times. We therefore do not count this operation toward the total FLOP count.
We also note that a quaternion operation is, on average, 16 times more costly than a real-valued operation. Hence, computing $\bm{K}_{2}^{\mathsf{H}}\bm{\nu}_{\mathrm{AA}}$ requires $16N_{A}N_{\mathrm{T}}\left(N_{A}(N_{A}-1)-1\right)$ FLOPs.
For the vector quaternion norm $\|\bm{\nu}_{\mathrm{AA}}\|^{2}$, we require 7 FLOPs (4 multiplications and 3 additions) per quaternion norm. Summing these values over all entries requires additional $N_{A}(N_{A}-1)/2-1$ scalar additions, for a total of $8\left(N_{A}(N_{A}-1)/2\right)-1$ FLOPs. The ratio then requires $N_{A}N_{\mathrm{T}}$ scalar–quaternion divisions, which cost $4N_{A}N_{\mathrm{T}}$ FLOPs. The subtraction corresponds to $N_{A}N_{\mathrm{T}}$ quaternion subtractions, which again cost $4N_{A}N_{\mathrm{T}}$ FLOPs. Thus, the expression in parentheses costs
[…]
$=16N_{A}N_{\mathrm{T}}\left(N_{A}(N_{A}-1)-1\right)+8\left(N_{A}(N_{A}-1)/2\right)-1$
$+8N_{A}N_{\mathrm{T}}.$ (79)
The factor $\bm{B}_{\mathrm{AT}}^{\mathsf{T}}/N_{\mathrm{A}}$ can be implemented by substituting the ones in $\bm{B}_{\mathrm{AT}}^{\mathsf{T}}$ with $1/N_{A}$, which we omit from the FLOP calculation. The remaining multiplication can be treated as a sparse multiplication of a real matrix with a quaternion vector, where each multiplication and addition costs 4 FLOPs. If implemented efficiently, computing one entry of $\bm{\hat{\chi}}_{\mathrm{T}}$ requires $N_{A}$ multiplications and $N_{A}-1$ additions, for a total of $8N_{A}-4$ FLOPs. This brings the total FLOP count of QD-MRC-SMDS to
FLOP s for QD-MRC-SMDS (80)
$=12M^{2}+16N_{A}N_{\mathrm{T}}\left(N_{A}(N_{A}-1)-1\right)$
$+8\left(N_{A}(N_{A}-1)/2\right)-1+8N_{A}N_{\mathrm{T}}$
$+N_{\mathrm{T}}(8N_{A}-4).$ (81)
-A4 Iterative QD-MRC-SMDS
Figure 15: Reconsidering Scenario II with larger room dimensions and more TNs.
We first note that if we substitute $\bm{\nu}_{\mathrm{AT}}^{(\tau_{\mathrm{max}})}=\bm{\nu}_{\mathrm{AT}}^{(0)}$ in (52), where $\bm{\nu}_{\mathrm{AT}}^{(0)}$ is defined in (51), then (52) becomes equivalent to QD-MRC-SMDS, for which the FLOP count was derived in Appendix A--A3.
The additional computation is (50), which involves a quaternion matrix–vector multiplication costing $16N_{A}N_{\mathrm{T}}(2M-1)$ FLOPs. The denominator is a quaternion-vector norm, which (as noted above) costs 8M18M-1 FLOPs. Finally, the ratio between a vector and a scalar requires $4N_{A}N_{T}$ FLOPs. This yields the total FLOP count for iterative QD-MRC-SMDS:
FLOP s for iterative QD-MRC-SMDS (82)
$=12M^{2}+16N_{A}N_{\mathrm{T}}\left(N_{A}(N_{A}-1)-1\right)$
$+8(N_{A}(N_{A}-1)/2)-1+8N_{A}N_{\mathrm{T}}$
$+N_{\mathrm{T}}(8N_{A}-4)$
$+16N_{A}N_{T}(2M-1)+8M-1+4N_{A}N_{T}.$ (83)
A single iteration is sufficient for convergence of iterative QD-MRC-SMDS; therefore, we do not include $\tau$ in the total FLOP count.
-B Effect of Simulation Environment on Localization Accuracy
In the main part of the article, the impact of changes in network topology (e.g., room size and number of TNs) on localization performance was not explicitly investigated. Additional experiments show that the main conclusion remains valid: QD-SMDS begins to outperform SMDS primarily when the angular measurements are highly unreliable.
To further support this claim, we revisit Scenario II in a new simulation environment with room dimensions of $100$ m (length) $\times$ $100$ m (width) $\times$ $30$ m (height). ANs are placed at five locations: the four upper corners of the room, specifically at $(x,y,z)=(0,0,10)$, $(100,0,30)$, $(100,100,30)$, and $(0,100,30)$, as well as the origin $(x,y,z)=(0,0,0)$. TNs are randomly deployed at 30 locations within the room interior, with their $\mathrm{x}$, $\mathrm{y}$, and $\mathrm{z}$ coordinates independently drawn from a uniform distribution. All other simulation settings, including the noise distributions, follow those in Section IV-C-1.
The results shown in Fig. 15 demonstrate that the conclusions presented in the main text remain unchanged: QD-SMDS outperforms SMDS only under sufficiently large angular errors. While the crossover point at which QD-SMDS surpasses SMDS may vary depending on the network topology, the qualitative performance advantage of QD-SMDS over SMDS remains consistent.